OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. 1. x^2 - 5x + 7 = 0 2. x^2 - 5x - 4 = 0 3. 16x^2 + 8x + 1 = 0 4. x^2 + 9x + 14 = 0 3 minutes agoEdit Question here are the options a.double root b.real and rational root c.real and irrational root d.imaginary root

OpenStudy (anonymous):

@mathstudent55 can you help?

OpenStudy (anonymous):

Do you know what to find the discriminant?

OpenStudy (anonymous):

no can you show me how to do 3 i figured out 1&2

OpenStudy (anonymous):

\[ 16x^2 + 8x + 1 = (4 x+1)^2=0\\ x=-\frac 14 \] double roots

OpenStudy (anonymous):

\[ b^2 - 4 a c = 64 - 4(16)1=0 \] double roots

OpenStudy (anonymous):

Do you understand it? I did it in two ways.

OpenStudy (anonymous):

what about #5?

OpenStudy (anonymous):

There is no number 5

OpenStudy (anonymous):

oh hah well what about this one? 2 = x^2 + 5x

OpenStudy (anonymous):

\[2 = x^2 + 5x\\ x^2 + 5 x -2=0\\ b^2-4ac= 25 +8=33 \\ \text { 33 is not a perfect square. Hence roots are irrationals } \]

OpenStudy (anonymous):

i got a few more...what about this one? 2m^2 + 3 = m

OpenStudy (anonymous):

I think, you should ask each question as a new problem, to avoid confusion for other students.

OpenStudy (anonymous):

Okay thanks anyway

OpenStudy (anonymous):

YW